OpenStudy (anonymous):

Rewrite the following Trigonimic functions as indicated: (use angles between 0 and 90 degrees) sin 150=sin____. sin 150=cos____. csc 45=sec____.

OpenStudy (dumbcow):

sin(x) = sin(180 - x) sin(x) = cos(x - 90)

OpenStudy (anonymous):

have you got to bring sex into it ishaan - lol

OpenStudy (anonymous):

lol :D

OpenStudy (anonymous):

sorry typo sorry ...sorry again sin 30 cos 60 sec45 actually x and c are adjacent keys sorry

OpenStudy (anonymous):

What about cos-30=cos___. cos-30=sin___. cot75=tan___.

OpenStudy (anonymous):

cos 150 sin 270 -60 =sin210 tan15

OpenStudy (dumbcow):

or cos(-30) = cos(30) cos(-30) = sin(60) cot(75) = tan(15)